### v2

# AnomalyGFM: Graph Foundation Model for Zero/Few-shot Anomaly Detection

Abstract.
Graph anomaly detection (GAD) aims to identify abnormal nodes that differ from the majority of the nodes in a graph, which has been attracting significant attention in recent years. Existing generalist graph models have achieved remarkable success in different graph tasks but struggle to generalize to the GAD task. This limitation arises from their difficulty in learning generalized knowledge for capturing the inherently infrequent, irregular and heterogeneous abnormality patterns in graphs from different domains. To address this challenge, we propose AnomalyGFM, a GAD-oriented graph foundation model that supports zero-shot inference and few-shot prompt tuning for GAD in diverse graph datasets. One key insight is that graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighbors). The residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistent way. This provides a driving force for the learning of graph-agnostic, discriminative prototypes for the normal and abnormal classes, which can be used to enable zero-shot GAD on new graphs, including very large-scale graphs. If there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support prompt tuning to leverage these nodes for better adaptation. Comprehensive experiments on 11 widely-used GAD datasets with real anomalies, covering social networks, finance networks, and co-review networks, demonstrate that AnomalyGFM significantly outperforms state-of-the-art competing methods under both zero- and few-shot GAD settings. Code is available at https://github.com/mala-lab/AnomalyGFM.
Anomaly Detection, Graph Neural Networks, Graph Anomaly Detection, Graph Foundation Models
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736843isbn: 979-8-4007-1454-2/2025/08conference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadaccs: Computing methodologies Machine learning
1. Introduction
Graph anomaly detection (GAD) aims to identify anomalous nodes within a graph that deviate significantly from others. It has found broad applications in areas such as financial fraud detection, transaction analysis, and social networks (Akoglu et al., 2015; Ma et al., 2021; Pang et al., 2021; Qiao et al., 2024). Most existing GAD methods, either supervised or unsupervised methods, follow a paradigm of training and inference on the same graph, making the trained models struggle to generalize to new/unseen graphs due to the distribution shift between the training and testing graphs (Qiao and Pang, 2023; Liu et al., 2021b; Ding et al., 2019). Further, this type of methods can become inapplicable in applications where graph data is not accessible during training due to data privacy or other data access issues. Generalist/foundation models for graphs have achieved remarkable progress in tackling these challenges in general tasks such as node classification, graph classification, and link prediction (Sun et al., 2023; Li et al., 2024; Zhao et al., 2024; Zi et al., 2024; Fang et al., 2024), but they are difficult to generalize to the GAD task. The main reason is that the generalized knowledge learned in these generalist models through generic pre-training or tuning methods inclines to the prevalent regular patterns, which differ significantly from the inherently infrequent, irregular, and heterogeneous abnormality patterns across the graphs (Qiao et al., 2024; Liu et al., 2024; Lin et al., 2024).
Figure 1. Given a set of graph datasets consisting of normal nodes (blue) and abnormal nodes (red) as in (a), AnomalyGFM aims to learn graph-agnostic prototypes for normal and abnormal nodes in a projected residual feature space. Compared to the original feature space where normal and abnormal patterns are irregularly distributed and differ significantly in different datasets as in (b), the residual feature space learned by AnomalyGFM in (c) is universally effective in separating anomalies from normal nodes in different graphs. In (b) and (c), AnomalyGFM is trained on Facebook (Xu et al., 2022) and evaluated directly on Amazon (Dou et al., 2020) and Question (Platonov et al., 2023); both feature spaces are learned by AnomalyGFM.
Very recently, there have been some efforts, such as ARC (Liu et al., 2024) and UNPrompt (Niu et al., 2024), dedicating to generalist models for GAD. Despite the effectiveness on some datasets, these methods fail to provide a universal GAD model that versatilely supports zero-shot inference and sample-efficient tuning in different application scenarios. For instance, UNPrompt (Niu et al., 2024) is designed exclusively for zero-shot inference, while ARC (Liu et al., 2024) is limited to few-shot tuning-based GAD within the same domain. Therefore, a foundational model being capable of effectively capturing abnormalities across graphs from different domains and supporting both zero-shot and few-shot GAD is still lacking.
To fill this gap, in this paper, we introduce AnomalyGFM, a GAD-oriented graph foundation model (GFM) that is capable of addressing both zero-shot and few-shot scenarios effectively for GAD. One key insight is that to avoid overfitting the pre-training graph data, graph-agnostic representations for normal and abnormal classes are required to support effective zero/few-shot GAD across different graphs. Motivated by this, AnomalyGFM is pre-trained to align data-independent, learnable normal and abnormal class prototypes with node representation residuals (i.e., representation deviation of a node from its neighboring nodes). Compared to the original feature space where normal and abnormal patterns are irregularly and heterogeneously distributed in different datasets (see Fig. 1 (b)), the residual features essentially project the node information into a unified feature space where we can effectively measure the abnormality of nodes from different graphs in a consistently identical way, as shown in Fig. 1 (c). This is because normal nodes, regardless of which graphs they are from, are expected to have small residual/deviation from its neighboring nodes, whereas such residuals for abnormal nodes are large (Qiao and Pang, 2023; Liu et al., 2024). By learning to align such residual features with the learnable normal and abnormal class prototypes, AnomalyGFM distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating strong generalization to GAD in new graphs without any further tuning/training. Furthermore, if there are few-shot labeled normal nodes available in the new graphs, AnomalyGFM can further support graph prompt tuning to leverage these nodes for better adaptation.
To summarize, the contributions of AnomalyGFM are three-fold:
  • We propose a new paradigm for GAD, aiming to pre-train a GAD-oriented GFM that achieves strong zero-shot generalization and supports sample-efficient graph prompt tuning for the GAD task.
  • We argue that learning graph-agnostic representations for normal and abnormal classes is a key ingredient of a GAD-oriented GFM. To justify this, we propose AnomalyGFM, the very first GAD-oriented GFM with strong zero-shot and few-shot generalization abilities. It is pre-trained to learn discriminative, graph-agnostic class prototypes with normal and abnormal residual features, and it supports few-shot graph prompt tuning for better adaptation.
  • A comprehensive benchmark on both zero-shot and few-shot settings using 11 real-world GAD datasets is established, on which i) AnomalyGFM performs significantly better than the state-of-the-art unsupervised, supervised, and generalist GAD methods, and ii) AnomalyGFM can scale up to very large graphs.
2. Related Work
2.1. Graph Anomaly Detection
Existing GAD methods can be generally classified into unsupervised and supervised methods (Ma et al., 2021; Qiao et al., 2024), depending on the assumption regarding the presence of normal and anomaly labels. The fully supervised approaches have recently achieved remarkable progress by leveraging both labeled normal and abnormal nodes (Tang et al., 2022; Liu et al., 2021a; Gao et al., 2023; Wang et al., 2023). These methods are primarily designed to improve message aggregation in graph neural networks (GNNs) by reducing heterophilic edges from both spectral and spatial perspectives, effectively mitigating the over-smoothing problem in GAD (Tang et al., 2023; Qiao et al., 2024). Fully supervised methods are effective in GAD as they treat the task as an imbalanced classification problem. However, the requirement for both normal and abnormal labels significantly hinders their applications to scenarios where labeled nodes are difficult to obtain.
Unsupervised methods, which typically assume the absence of both normal and abnormal labels, have garnered significant attention due to their more practical setting assumption on data labels. They generally incorporate some conventional techniques such as reconstruction (Ding et al., 2019), one-class classification (Wang et al., 2021; Zhou et al., 2021; Qiao and Pang, 2023), contrastive learning (Liu et al., 2021b; Pan et al., 2023), and adversarial learning (Ding et al., 2021) into graph learning to capture the normal patterns within the graph and then assign an anomaly score for each node based on its deviation from the normal patterns. However, these methods still follow the paradigm of training and inference on the same graph, making them struggle to generalize to new/unseen graphs due to the distribution shift between training and testing set.
2.2. Foundation Models for GAD
Generalist models have recently achieved significant progress on non-graph data by leveraging large pre-trained models to facilitate generalized pattern recognition in diverse downstream tasks, such as generalist anomaly detection on images (Zhou et al., 2023; Zhu and Pang, 2024). However, applying these methods to graph data remains highly challenging due to the absence of such pre-trained models (Sun et al., 2023; Liu et al., 2023; Wen et al., 2023) on graph data. The main challenge in designing a GFM is capturing invariance across diverse graphs while mapping them into a shared representation space to enable positive transfer between training and inference (Mao et al., [n. d.]; Beaini et al., 2023; Galkin et al., 2023). Currently, most GFM models use prompt learning to enable the knowledge transfer across graphs for general graph tasks (Liu et al., 2023; Fang et al., 2024). However, they still struggle to generalize to the GAD task due to the inherently infrequent, irregular, and heterogeneous abnormality patterns in GAD datasets (Liu et al., 2024; Niu et al., 2024; Qiao et al., 2024). Therefore, some recent efforts attempt to devise the GFMs for GAD (Niu et al., 2024; Liu et al., 2024). ARC (Liu et al., 2024) is one of such methods, a fine-tuning GFMs for GAD, enabling a “one-for-all” GAD model through a fine-tuning mechanism. UNPrompt (Niu et al., 2024) is the first zero-shot generalist GAD method that learns generalized neighborhood prompts, allowing latent node attribute predictability to serve as a generalized anomaly measure. Different from ARC and UNPrompt, which address the GAD-oriented GFMs partially under few-shot or zero-shot settings, AnomalyGFM can effectively capture abnormalities across graphs from different domains while supporting both zero-shot inference and few-shot prompt fine-tuning/inference. This offers more versatile abilities for a GFM-based GAD approach.
3. The Proposed AnomalyGFM
3.1. Preliminaries
Notation and Conventional GAD. An attributed graph is denoted by $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ represents the node set and $\mathcal{E}$ represents the edge set. The graph can also be denoted as $\mathcal{G}=(\mathbf{A},\mathbf{X})$ where $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right]^{T}\in\mathbb{R}^{N\times\tilde{d}}$ is the attribute of the nodes and $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacent matrix of the graph with $\mathbf{A}_{ij}=0$ representing there is no edge between node $v_{i}$ and node $v_{j}$ and $A_{ij}=1$ otherwise. In GAD, we denote the normal node set as ${\mathcal{V}}_{n}$ and abnormal node set ${\mathcal{V}}_{a}$, corresponding to the anomaly labels ${\bf{y}}\in{\{0,1\}^{N}}$ with ($y_{i}=0$ denotes a normal node $v_{i}$ and $y_{i}=1$ denotes abnormal). Typically, the number of normal nodes is significantly larger than the abnormal nodes in the GAD datasets, i.e., $\left|{{{\mathcal{V}}_{n}}}\right|\gg\left|{{{\mathcal{V}}_{a}}}\right|$. The conventional GAD is formulated as learning a scoring function $f:\mathcal{G}\rightarrow\mathbb{R}$ and making inference on the same $\mathcal{G}$, in which $f$ is expected to assign every node an anomaly score such that $f(v)<f(v^{\prime})$, $\forall v\in\mathcal{V}_{n},v^{\prime}\in\mathcal{V}_{a}$.
Zero-shot GAD. Different from the conventional GAD that trains and inferences on the same graph, zero-shot GAD aims to train a generalist model on $\mathcal{G}_{Train}$ and then apply the learned model to the new test graphs $\mathcal{G}_{Test}$ without any further tuning/retraining.
Few-shot GAD. In the few-shot GAD setting, the model is fine-tuned on $\mathcal{G}_{Test}$ with small labeled nodes and then applied to the rest of the unlabelled nodes in the test graph $\mathcal{G}_{Test}$. We denote the labeled node in the test graph as $\mathcal{V}_{l}^{Test}$ while the unlabeled nodes in the test graph are denoted as $\mathcal{V}_{u}^{Test}=\mathcal{V}^{Test}\setminus\mathcal{V}_{l}^{Test}$.
Feature Unification. Due to the feature dimension difference across the graphs, we need to align the node features/attributes into a shared feature space to ease the feature heterogeneity among the graphs. Following the previous studies (Liu et al., 2024; Niu et al., 2024), we employ singular value decomposition (SVD), a simple, efficient dimensionality reduction technique that can approximately preserve the distance relationship of the data by projecting high-dimensional data into a low-dimensional subspace for unification, For each dataset $\mathbf{X}^{(i)}$ with dimension $d^{(i)}$, it can be formulated as:
[…]
where $d^{\prime}$ is the common dimensionalities and $N^{(i)}$ is the number of node for graph $\mathcal{G}_{i}$
GNN for Representation Learning. Due to its simplicity and effectiveness, GNN is employed to learn the representation of each graph which can be formulated as follows,
(2) ${\mathbf{H}^{(\ell)}=\operatorname{GNN}\left(\mathbf{A},\mathbf{H}^{(\ell-1)};\mathbf{W}^{(\ell)}\right),}$
where ${\bf{H}}^{(\ell)}\in\mathbb{R}^{N\times h^{(l)}}$, ${\bf{H}}^{(\ell-1)}\in\mathbb{R}^{N\times h^{(l-1)}}$ are the representations of all $N$ nodes in the ${(\ell)}$-th layer and ${(\ell-1)}$-th layer, respectively, $h^{(l)}$ is the dimension of ${(\ell)}$-th layer, ${\bf{W}}^{(\ell)}$ are the learnable parameters, and ${\bf{H}}^{(0)}=\tilde{\mathbf{X}}$. ${{{\bf{H}}}=\left\{{{{\bf{h}}_{1}},{{\bf{h}}_{2}},\ldots,{{\bf{h}}_{N}}}\right\}}$ is a set of representations of $N$ nodes in the last GNN layer. In this paper, we adopt a 2-layer GCN to model the graph due to its efficient and simple architecture.
3.2. Approach Overview
AnomalyGFM aims to distill discriminative node residual features in a unified feature space into two graph-agnostic prototypes for the normal and abnormal classes, facilitating the establishment of a foundation model that supports effective zero/few-shot GAD across graphs. Specifically, AnomalyGFM consists of pre-training based on the prototype alignment, few-shot graph prompt fine-tuning, and zero-shot/few-shot inference, as shown in Fig. 2. In the pre-training step, we aim to align the data-independent and learnable normal class and abnormal class prototypes with the nodes’ residual features to achieve the distillation of graph-agnostic discriminative class-level representations. In the few-shot graph prompt fine-tuning, the labeled normal nodes’ residual features are utilized to guide a better adaptation of the normal class prototype. In the inference, we devise zero-shot inference on new graphs, including subgraph-based inference for very large-scale graphs, and few-shot inference, both of which compute anomaly scores by measuring the similarity between the residual features of nodes and the two class prototypes.
Figure 2. Overview of AnomalyGFM. (a) During the pre-training, AnomalyGFM aims to align two learnable class-level prototypes with the representation residuals of normal and abnormal nodes to learn discriminative graph-agnostic prototypes on an auxiliary labeled graph. (b) During inference, the anomaly score is determined in a consistent way using similarity between node representation residual and prototypes where all learnable parameters are frozen. Given a target graph with few-shot labeled normal nodes, these nodes are used to better adapt the pre-trained normal class prototypes to the target graph through a one-class prompt tuning method.
3.3. Pre-training via Prototype Alignment in a Unified Feature Space
Node Representation Residual. Node representation residual can be seen as the deviation at the representation level between the target node and its neighboring nodes. The deviation in the representation between the target node and its neighbors is expected to be small for normal nodes, regardless of the graph domain, while the deviation for abnormal nodes is relatively large. Therefore, the node representation residual offers a consistent method to effectively capture heterogeneous abnormalities across graphs. Assume the node representation outputted by the GNNs as $\mathbf{h}_{i}\in\mathbb{R}^{d}$, where $d$ is the dimension of representation, the residual feature $\mathbf{r}_{i}\in\mathbb{R}^{d}$ for the node $v_{i}$ is formally defined as follows:
(3) ${\mathbf{r}_{i}=\mathbf{h}_{i}-\frac{1}{\left|\mathcal{N}\left(v_{i}\right)\right|}\sum_{v_{j}\in\mathcal{N}\left(v_{i}\right)}\mathbf{h}_{j},}$
where $\mathcal{N}\left(v_{i}\right)$ represents the neighbor set of node $v_{i}$. As a graph-agnostic representation, the node representation residual will be assigned to each node as its new feature in the unified feature space.
Graph-agnostic Prototype Alignment. To exploit the anomaly discriminability in node representation residuals, we learn two data-independent learnable prototypes for the two classes that align with the normal and abnormal residual features, respectively. The alignment essentially distills the residual feature-based anomaly discriminability into the two graph-agnostic prototypes, facilitating a strong generalization to GAD in new graphs. Formally, we integrate two learnable mapping functions ${\bf p}_{n}=\Phi({\bf{z}}_{n};\Theta_{n})$, where ${\bf{z}}_{n},{\bf p}_{n}\in\mathbb{R}^{T}$ and ${\bf p}_{a}=\Phi({\bf{z}}_{a};\Theta_{a})$, where ${\bf{z}}_{a},{\bf p}_{a}\in\mathbb{R}^{T}$ to map their initializations (e.g., prototypes initialized with vectors drawn from a prior distribution such as Gaussian distribution) into the prototype space to serve as normal class prototype $\mathbf{p}_{n}$ and abnormal class prototype $\mathbf{p}_{a}$, respectively. Both $\Theta_{n}$ and $\Theta_{a}$ are the learnable parameters, and $T$ is the dimension of the learnable prototype, which is set to match the dimensionality of the node representation $d$. The Gaussian prior distributions ${\bf{z}}_{n}$ and ${\bf{z}}_{a}$ are uniformly defined as $\mathcal{N}(\mu,\sigma)$ for the normal and abnormal classes, where ${\mu}$ is the mean of the distribution and $\sigma$ is its covariance.
The pre-training of AnomalyGFM is performed by aligning the two learnable prototypes with the abnormal residual features and normal residual features respectively using an alignment loss function $L_{Alignment}$ defined as follows:
(4) ${{L_{Alignment{\rm{}}}}=\sum\limits_{i=1}^{\left|{{{\mathcal{V}}}}\right|}{{I_{{y_{i=0}}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{n}}}\right\|_{2}^{2}+{I_{{y_{i=1}}}}\left\|{{{\bf{r}}_{i}}-{{\bf{p}}_{a}}}\right\|_{2}^{2}}\,,$
where ${I_{{y_{i=0}}}}$ is an indicator function, which takes the value of 1 when the condition is satisfied. This distills the graph-agnostic discriminability in the residual feature space into the two prototypes. To enhance the anomaly discriminability in the node representation residual, we utilize a standard binary classification loss $L_{BCE}$ to learn discriminative representations for the normal and abnormal nodes in the original node representation space:
(5) ${L_{BCE}}=\sum\limits_{i=1}^{\left|{{{\mathcal{V}}}}\right|}{{y_{i}}}\log\left({{p_{i}}}\right)+\left({1-{y_{i}}}\right)\log\left({1-{p_{i}}}\right),$
where ${{p_{i}}={f_{\theta}}\left({{\bf h}_{i}}\right)}$ and $f_{\theta}(\cdot)$ is a Multi-Layer Perceptron (MLP) classifier with training parameters $\theta$, which maps the representation to the probability for a node being an anomaly. Finally, we jointly optimize the ${L_{Alignment}}$ and $L_{BCE}$ via the following total loss :
(6) ${{L_{total{\rm{}}}}={L_{BCE}}+\alpha{L_{Alignment{\rm{}}}}},$
where $\alpha$ is a hyperparameter to control the importance of the proposed alignment module. Note that the pre-training is performed on auxiliary graph datasets rather than target graph datasets.
3.4. Few-shot Prompt Tuning
To enable the few-shot prompt tuning in the scenarios where few-shot labeled normal nodes are available in the test/target graph, we utilize these limited nodes to have a prompt tuning of AnomalyGFM in the prototype learning module while keeping the GNN parameters frozen.
Normal Prototype-based Prompting. To fully leverage the labeled normal nodes during the fine-tuning step, we introduce a small learnable prompt into the normal class prototype to better align it with the normal node representation residuals on the test graph during a prompt fine-tuning step. In addition, we also add an adaptation layer that offers more learning flexibility to the tuning. They are designed to refine the pre-trained normal class prototype ${\bf{p}}_{n}$ to a new normal prototype $\bf{p}_{n}^{\prime}$, ensuring a better alignment with the normal class residual feature in the target graphs. Formally, the new normal prototype ${\bf{p}^{\prime}_{n}}$ is defined as:
(7) ${\bf{p}^{\prime}_{n}}={{\bf{p}}_{n}}+{g}({{\bf{p}}_{n};\phi})+\Psi_{\mathbf{p}_{n}},$
where $\Psi_{\mathbf{p}_{n}}\in\mathbb{R}^{d}$ is the learnable prompt while ${g}({{\bf{p}}_{n};\phi})$ is the added adaptation layer with parameters $\phi$, both of which are of the same dimensionality as ${{\bf{p}}_{n}}$. The learned prompts and the adapted prototype complements the pre-trained normal prototype in the target graph, so we fuse these three vectors to obtain the updated normal class prototype ${\bf{p}^{\prime}_{n}}$.
One-class Prompt Tuning. During the prompt tuning, the original learnable parameters of AnomalyGFM are all frozen. The few-shot labeled normal nodes are solely utilized to learn the new parameters $\phi$ and $\Psi$ for refining the normal class prototype. Specifically, given some labeled normal node $\mathcal{V}_{l}^{Test}$ in the test graph, we first calculate the residual feature ${\bf{r}}_{i}$ of the labeled normal node $v_{i}\in\mathcal{V}_{l}^{Test}$ using Eq. 3. Then the graph prompt is optimized using the following one-class tuning loss $L_{{{pt}}}$, aiming to minimize the Euclidean distance between the residual features of labeled normal node ${\bf r}_{i}$ and the new trainable normal class prototype $\bf{p}^{\prime}_{n}$ composed by the prompt and the pre-trained normal prototype.
(8) ${L_{{{pt}}}}=\sum\limits_{i=1}^{\left|{{\mathcal{V}}_{l}^{Test}}\right|}\left\|{{{\bf{r}}_{i}}-{\bf{p}^{\prime}_{n}}}\right\|_{2}^{2},$
where the residual feature ${\bf r}_{i}$ is fixed since the learnable parameters in GNN are frozen while ${\bf p}^{\prime}_{n}$ is adjusted based on the newly added trainable parameters.
3.5. GAD Using AnomalyGFM
Zero-shot Inference. To build a foundation model for GAD with zero-shot learning capabilities, we directly apply AnomalyGFM pre-trained on auxiliary graphs to the test graph $\mathcal{G}_{test}$, meaning that $\Theta_{n}$ and $\Theta_{a}$ in $\Phi({\bf{z}}_{n};\Theta_{n})$ and $\Phi({\bf{z}}_{a};\Theta_{a})$ and the trainable parameters $\bf{W}$ in GNN are all frozen. This process allows us to obtain the residual features for each node in the test graph, the graph-agonistic normal prototype ${\bf{p}}_{n}$ and abnormal prototype ${\bf{p}}_{a}$. The anomaly score is determined by computing the similarity of the target node’s representations residual with the two prototypes ${\bf{p}}_{a}$ and ${\bf{p}}_{n}$. Formally, the anomaly score $s_{i}$ for the node $v_{i}$ is defined as,
(9) ${{s_{i}}=\exp\left({{\bf{r}}_{i}^{\rm{T}}{{\bf{p}}_{a}}}\right)+\beta\exp\left(-{{\bf{r}}_{i}^{\rm{T}}{{\bf{p}}_{n}}}\right)},$
where $\beta$ is the hyperparameter that controls the weight of two parts. Note that the anomaly scoring calculation consists of the direct similarity with ${\bf p}_{a}$ and the inverse similarity with ${\bf p}_{n}$.
Few-shot Inference. In the few-shot inference, the original parameters in GNN, mapping functions, and newly added parameters for refining the normal class prototype are frozen. The anomaly score of ${{\mathcal{V}}_{u}^{Test}}$ is calculated based on the prompt-tuning-based updated normal class prototype $\bf{p}_{n}^{\prime}$ and the pre-trained abnormal class prototype ${\bf{p}}_{a}$, which is defined as:
(10) ${{s_{i}}=\exp\left({{\bf{r}}_{i}^{\rm{T}}{{\bf{p}}_{a}}}\right)+\beta\exp\left(-{{\bf{r}}_{i}^{\rm{T}}\bf{p}_{n}^{\prime}}\right)}.$
Inference on Very Large-scale Graphs. Most existing GAD methods typically require loading all nodes of the graph, which often leads to poor scalability on large graphs during inference. By learning the graph-agnostic prototypes, AnomalyGFM can generalize to very large-scale graphs through a subgraph-based inference approach. AnonalyGFM can effectively infer the anomaly score without considering the entire graph structure, eliminating the bottleneck of loading the full graph for GAD inference. This subgraph inference is also desired in privacy-sensitive settings where we do not want to reveal the entire graph structure to the detection models. Specifically, AnomalyGFM operates in a way that given a test node $v_{i}$ and its subgraph $\mathcal{S}\left(v_{i}\right)$ with size $s$ are given, it first computes its residual feature ${\bf{r}}_{i}$ as:
(11) ${\mathbf{r}_{i}=\mathbf{h}_{i}-\frac{1}{\left|\mathcal{S}\left(v_{i}\right)\right|}\sum_{v_{j}\in\mathcal{S}\left(v_{i}\right)}\mathbf{h}_{j}}\,,$
where ${\mathcal{S}}\left(v_{i}\right)$ is extracted using a random walk starting with the target node $v_{i}$. Then, AnomalyGFM infers the anomaly score using the similarity of the residual feature to the normal and abnormal class prototypes as defined in Eq. (9).
4. Experiments
4.1. Experimental Setup
Table 1. Key statistics of GAD datasets used in our experiments. Ano. indicates the number (rate) of anomalies in the graph and Sim. indicates the global average edge similarity.
Dataset Domain #Nodes #Edges #Feat. Ano. Sim.
Facebook Social Networks 1,081 55,104 576 25(2.31%) 0.690
Reddit Social Networks 10,984 168,016 64 366(3.33%) 0.997
Amazon Co-review 10,244 175,608 25 693(6.66%) 0.645
Disney Co-purchase 124 335 28 6(4.8%) 0.804
Amazon-all Co-review 11,944 4,398,392 25 821(6.87%) 0.645
YelpChi-all Co-review 45,941 3,846,979 32 6,674(14.52%) 0.905
Tolokers Work Collaboration 11,758 519,000 10 2,566(21.8%) 0.814
Question Social Networks 48,921 153,540 301 1,460(2.98%) 0.679
Elliptic Bitcoin Transaction 46,564 73,248 93 4,545 (9.8%) 0.356
T-Finance Transaction Record 39,357 21,222,543 10 1,803(4.6%) 0.107
T-Social Social Friendship 5,781,065 73,105,508 10 174,280(3.0%) 0.307
Datasets. We utilize 11 benchmark datasets from social networks, finance, and co-review domains for evaluation. These datasets are summarized in Table 1, with additional details provided in Appendix A. The social network datasets consist of Facebook (Xu et al., 2022), Reddit (Kumar et al., 2019), Question (Platonov et al., 2023) and T-Social (Tang et al., 2022). The financial datasets include T-Finance (Tang et al., 2022) and Elliptic (Weber et al., 2019). For the co-review category, we use Amazon (Dou et al., 2020), Amazon-all (Dou et al., 2020), and YelpChi-all (Dou et al., 2020). Additionally, the Disney (Sánchez et al., 2013) and Tolokers (Platonov et al., 2023) datasets, derived from co-purchase and work collaboration networks, are also included.
Table 2. AUROC and AUPRC results for zero-shot GAD on nine real-world datasets, with the models trained on Facebook only. For each dataset, the best performance per column within each metric is boldfaced, with the second-best underlined. Avg. denotes the average performance. P-values are from one-tailed Wilcoxon signed rank tests.
Metric Method Dataset Avg. p-value
Reddit Amazon Disney Aamzon-all YelpChi-all Tolokers Question Elliptic T-Finance
AUROC Unsupervised Methods
AnomalyDAE (ICASSP’20) 0.5016 0.5818 0.4853 0.7228 0.5002 0.5948 0.4311 0.4197 0.2324 0.4966 0.007
CoLA (TNNLS’21) 0.4623 0.4580 0.4696 0.4091 0.4879 0.4501 0.4945 0.5572 0.4889 0.4752 0.003
TAM (NeurIPS’23) 0.5725 0.4720 0.4773 0.7543 0.4216 0.5351 0.5119 0.3282 0.2990 0.4857 0.003
GADAM (ICLR’24) 0.4532 0.6646 0.4288 0.5959 0.4829 0.4832 0.5594 0.3922 0.1382 0.4664 0.007
Supervised Methods
GCN (ICLR’17) 0.5645 0.5988 0.5000 0.7195 0.5486 0.5319 0.5161 0.7640 0.2345 0.5531 0.039
GAT (ICLR’18) 0.5000 0.4981 0.5175 0.5005 0.4802 0.5030 0.4577 0.6588 0.5072 0.5136 0.007
BWGNN (ICML’22) 0.5208 0.4769 0.6073 0.3648 0.5282 0.4877 0.4404 0.5843 0.5457 0.5062 0.003
GHRN (WebConf’23) 0.5253 0.4560 0.5336 0.3382 0.5125 0.4860 0.4535 0.5400 0.5324 0.4863 0.003
XGBGraph (NeurIPS’23) 0.4601 0.4179 0.6692 0.7950 0.4945 0.5462 0.5095 0.4274 0.3402 0.5177 0.003
Generalist Methods
GraphPrompt (WebConf’23) 0.4677 0.4904 0.5192 0.3215 0.4976 0.4779 0.4204 0.3221 0.5405 0.4508 0.003
UNPrompt (Arxiv’24) 0.5337 0.7525 0.6412 0.7962 0.5558 0.6853 0.4757 0.5901 0.2318 0.5847 0.074
AnomalyGFM 0.5974 0.8417 0.6751 0.9032 0.5791 0.5843 0.5280 0.6195 0.5614 0.6544 /
AUPRC Unsupervised Methods
AnomalyDAE (ICASSP’20) 0.0327 0.0833 0.0566 0.1921 0.1484 0.1876 0.0241 0.0798 0.0274 0.0924 0.003
CoLA (TNNLS’21) 0.0391 0.0669 0.0701 0.0861 0.1466 0.0848 0.0292 0.0998 0.0430 0.0739 0.007
TAM (NeurIPS’23) 0.0413 0.0666 0.0628 0.1736 0.1240 0.0970 0.0307 0.0697 0.0332 0.0776 0.007
GADAM (ICLR’24) 0.0293 0.1562 0.0651 0.1595 0.1371 0.1001 0.0395 0.0733 0.0261 0.0873 0.003
Supervised Methods
GCN (ICLR’17) 0.0439 0.0891 0.0484 0.1536 0.1735 0.1060 0.0387 0.1963 0.0274 0.0974 0.074
GAT (ICLR’18) 0.0329 0.0688 0.0530 0.0696 0.1400 0.0822 0.0259 0.1366 0.0463 0.0728 0.003
BWGNN (ICML’22) 0.0389 0.0652 0.0624 0.0586 0.1605 0.1030 0.0257 0.1158 0.0479 0.0753 0.007
GHRN (WebConf’23) 0.0407 0.0633 0.0519 0.0569 0.1505 0.0957 0.0259 0.1148 0.0457 0.0717 0.007
XGBGraph (NeurIPS’23) 0.0330 0.0536 0.1215 0.2307 0.1449 0.1256 0.0306 0.0816 0.0754 0.0996 0.027
Generalist Methods
GraphPrompt (WebConf’23) 0.0334 0.0661 0.0610 0.0666 0.1542 0.2070 0.0266 0.0664 0.0492 0.0811 0.003
UNPrompt (Arxiv’24) 0.0351 0.1602 0.1236 0.2430 0.1810 0.2219 0.0348 0.1278 0.0279 0.1283 0.003
AnomalyGFM 0.0387 0.5790 0.1242 0.6820 0.1819 0.2749 0.0397 0.1371 0.0593 0.2352 /
Zero-shot Competing Methods. For the zero-shot setting, we evaluate AnomalyGFM against the state-of-the-art approaches from three main categories as follows (1) unsupervised methods: AnomalyDAE (Fan et al., 2020), CoLA (Liu et al., 2021b), TAM (Qiao and Pang, 2023), and GADAM (Chen et al., 2024); (2) supervised methods: GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), BWGNN (Tang et al., 2022), GHRN (Gao et al., 2023), and XGBGraph (Tang et al., 2023); and (3) general GFM methods, including GraphPrompt (Liu et al., 2023) for general graph tasks, and zero-shot GAD UNPrompt (Niu et al., 2024).
Few-shot Competing Methods. In few-shot setting, three generalist methods are icluded, i.e.GPPT (Sun et al., 2022), GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024).
Evaluation Metrics. Following the previous studies (Liu et al., 2024; Tang et al., 2022; Qiao and Pang, 2023), two widely-used metrics, AUROC and AUPRC, are used to evaluate the performance of all methods. For both metrics, a higher value denotes better performance. Moreover, for each method, we report the average performance for 3 independent runs with different random seeds. One-tailed Wilcoxon signed rank test is performed to evaluate the statistical significance of the performance.
Implementation Details. For all competing methods and AnomalyGFM, the code is implemented with Pytorch 1.11 , PyG 2.1.0, DGL 1.0.1, and Python 3.8.19. All experiments are conducted on a Linux server with an Intel CPU (Intel Xeon Gold 6346 3.1GHz) and an Nvidia RTX3090 GPU with 24 GB GPU memory. The pre-train model is optimized using the Adam optimizer (Kingma, 2014) with 300 epochs and a learning rate of $41e-$ 4 by default. The dimension of hidden space for representation and the prototype size are uniformly set to 300. The parameter $\alpha$ in the loss function is set to one by default. The setting of $\beta$ is determined based on the prior we gain from the global average similarity information provided in Table 1 (see Sec. 6 for details). The $\mu$ and $\sigma$ of the Gaussian distribution for both normal and abnormal class prototypes are set to zero and one by default. The size of the subgraph in subgraph inference is set to five by default. The implementation of competing methods is based on their publicly available official source code.
Table 3. Results on the nine real-world GAD datasets under 1/5/10-shot with models tuned on the few-shot data.
Metric Setting Method Dataset Avg. p-value
Reddit Amazon Disney Aamzon-all YelpChi-all Tolokers Question Elliptic T-Finance
AUROC 1-shot GPPT (KDD’22) 0.5000 0.5303 0.4997 0.5010 0.5000 0.5061 0.4921 0.6162 0.3647 0.5011 0.003
GraphPrompt (WebConf’23) 0.4216 0.4882 0.4223 0.2631 0.4811 0.5328 0.4086 0.6001 0.4000 0.4464 0.004
ARC (NeurIPS’24) 0.4899 0.4571 0.3578 0.4570 0.4910 0.4667 0.5865 0.2904 0.2484 0.4272 0.008
AnomalyGFM 0.5922 0.8531 0.6649 0.8972 0.5872 0.5898 0.5303 0.6199 0.5916 0.6584 /
5-shot GPPT (KDD’22) 0.5000 0.5098 0.5000 0.5051 0.5000 0.5181 0.4959 0.5736 0.2609 0.4848 0.003
GraphPrompt (WebConf’23) 0.4406 0.4900 0.6497 0.4726 0.5359 0.5381 0.4069 0.6012 0.4069 0.5046 0.003
ARC (NeurIPS’24) 0.4720 0.4458 0.4435 0.4473 0.5112 0.4746 0.5906 0.2714 0.2168 0.4303 0.007
AnomalyGFM 0.6023 0.8600 0.6613 0.9011 0.5951 0.6095 0.5426 0.6119 0.6248 0.6676 /
10-shot GPPT (KDD’22) 0.5000 0.5087 0.4769 0.5023 0.5000 0.4971 0.5047 0.4212 0.5539 0.4961 0.003
GraphPrompt (WebConf’23) 0.4321 0.4906 0.6314 0.7167 0.5367 0.5329 0.3826 0.6221 0.4260 0.5301 0.007
ARC (NeurIPS’24) 0.4867 0.4323 0.4769 0.4467 0.5145 0.4786 0.5901 0.2644 0.2298 0.4355 0.003
AnomalyGFM 0.6252 0.8649 0.6649 0.9215 0.6064 0.6140 0.5611 0.6303 0.6283 0.6796 /
AUPRC 1-shot GPPT (KDD’22) 0.0333 0.0766 0.0488 0.0687 0.1453 0.2204 0.0294 0.1239 0.0432 0.0877 0.003
GraphPrompt (WebConf’23) 0.0283 0.0680 0.0486 0.0426 0.1113 0.2321 0.0448 0.1108 0.0302 0.0796 0.012
ARC (NeurIPS’24) 0.0332 0.0581 0.0453 0.0590 0.1402 0.2122 0.0468 0.0701 0.0277 0.0769 0.011
AnomalyGFM 0.0398 0.5801 0.1223 0.6921 0.1852 0.2786 0.0332 0.1401 0.0601 0.2368 /
5-shot GPPT (KDD’22) 0.0333 0.0692 0.0504 0.0716 0.1453 0.2265 0.0297 0.1127 0.0365 0.0861 0.004
GraphPrompt (WebConf’23) 0.0285 0.0681 0.0892 0.0600 0.1661 0.2957 0.0327 0.1416 0.0360 0.1019 0.009
ARC (NeurIPS’24) 0.0312 0.0571 0.0546 0.0572 0.1464 0.2150 0.0471 0.0726 0.0267 0.0786 0.011
AnomalyGFM 0.0401 0.5831 0.1257 0.6985 0.1918 0.2866 0.0336 0.1437 0.0622 0.2405 /
10-shot GPPT (KDD’22) 0.0334 0.0691 0.0526 0.0698 0.1453 0.2178 0.0301 0.0905 0.0511 0.0844 0.004
GraphPrompt (WebConf’23) 0.0278 0.0681 0.0848 0.1427 0.1649 0.2922 0.0263 0.1421 0.0382 0.1096 0.007
ARC (NeurIPS’24) 0.0327 0.0557 0.0743 0.0583 0.1491 0.2168 0.0463 0.0677 0.0272 0.0809 0.011
AnomalyGFM 0.0444 0.5895 0.1399 0.7124 0.1990 0.2897 0.0346 0.1570 0.0644 0.2478 /
4.2. Zero-shot GAD Performance
The AUROC and AUPRC results under the zero-shot setting are shown in Table 2 where the model is trained on Facebook and inference is made on other datasets without any further tuning. Similar results can be found when the models are trained on the Amazon dataset in Appendix B.2. From the table, we observe (1) AnomalyGFM consistently outperforms the three types of competing methods, including unsupervised methods, supervised methods, and generalist methods on most datasets. Specifically, AnomalyGFM achieves the best AUROC performance on six datasets and the best AUPRC performance on the five datasets, with up to 11% improvement in AUROC and 44% improvement in AUPRC when compared to the best-competing method. The main reason is that AnomalyGFM performs inference based on two discriminative graph-agonistic prototypes distilled from the node residual feature, supporting strong generalization across the datasets. (2) Unsupervised methods often fail to produce consistently good results across the datasets due to the fact that each of their model designs is devised in a way to focus on some specific abnormal patterns for the training graphs, making them struggle to generalize to the new graphs. (3) Most supervised methods perform well in zero-shot inference, particularly simple approaches like GCN, which achieves top performance on the Elliptic dataset. This can be attributed to their simple model design, avoiding heavy overfitting to the training graph. However, the performance still underperforms AnomalyGFM, demonstrating that the node representation residual is more generalized and our prototype alignment module provides a strong add-on to the binary loss function in our method. (4) Generalist methods, such as GraphPrompt, which are designed for general node/graph classification, struggle to perform well on GAD. This is due to the irregular nature of the abnormalities in different GAD datasets, hindering the applicability of the regular patterns/knowledge these generalist method learned from other tasks to GAD. AnomalyGFM outperforms UNPrompt on all datasets except Tolokers, which can be attributed to the node representation residual-based class prototypes offers more generalization power than the latent attribute predictability used in UNPrompt. Further, our statistical significance test results show that the superiority of AnomalyGFM over all competing methods is significant at the 90%/95% confidence level in both metrics.
4.3. Few-shot GAD Performance
To demonstrate the effectiveness in the few-shot scenarios, the comparison of AnomalyGFM to GraphPrompt (Liu et al., 2023) and ARC (Liu et al., 2024) is done under 1/5/10-shot. The AUROC and AUPRC results are shown in Table 3. In the few-shot GAD scenario, AnomalyGFM outperforms these two generalist methods on most datasets in both AUROC and AUPRC. GraphPrompt continues to underperform AnomalyGFM since the presence of only one class of samples renders its fine-tuning ineffective, leading to suboptimal performance. ARC also leverages residual features
Table 4. Subgraph inference on the very large-scale graphs. ‘/’ indicates that the model cannot handle the dataset.
Metric Method Dataset
T-Finance T-Social
AUROC Unsupervised Methods
TAM (NeurIPS’23) 0.2990 /
GADAM (ICLR’24) 0.1382 0.5155
Supervised Methods
BWGNN (ICML’22) 0.5457 0.4964
GHRN (WebConf’23) 0.5324 0.4934
XGBGraph (NeurIPS’23) 0.3402 0.4602
AnomalyGFM 0.7852 0.5991
AUPRC Unsupervised Methods
TAM (NeurIPS’23) 0.0332 /
GADAM (ICLR’24) 0.0261 0.0285
Supervised Methods
BWGNN (ICML’22) 0.0479 0.0301
GHRN (WebConf’23) 0.0457 0.0303
XGBGraph (NeurIPS’23) 0.0754 0.0305
AnomalyGFM 0.1059 0.0398
and is specifically designed for few-shot GAD datasets achieving the best performance on the Question and the second performance on most datasets. Compared to ARC, AnomalyGFM achieves better sample-efficiency tuning due to the learning of class-level prototypes that are agnostic to different domains of graphs, enhancing the generalization across the graph, thereby yielding the best performance on the eight datasets. Tuning the abnormal prototype is also possible when a few labeled anomalies are available and the results are provided in App. B.1.
4.4. Scale up to Very Large Graphs
We choose two large-scale graphs T-Finance (with a large number of edges) and T-Social (with a large number of nodes) to evaluate the generalization of AnomalyGFM on very large graphs. The results are shown in Table 4. We can observe that AnomalyGFM outperforms the unsupervised methods and supervised methods on these datasets. This is mainly because the learned data-independent prototypes can effectively measure the abnormality of node representation residual, even in the presence of subgraphs. The deviations of the target node from the contextual node are still preserved in the subgraph, thereby enabling effective subgraph-based inference.
Table 5. Ablation study results for AnomalyGFM and its variants.
Metric Method Dataset
Reddit Amazon Disney Amazon-all YelpChi-all Tolokers Question Elliptic T-Finance Avg.
AUROC FA 0.5697 0.5680 0.6073 0.5097 0.5211 0.5594 0.5326 0.7765 0.6674 0.5901
BCE 0.5445 0.4308 0.7542 0.5545 0.5232 0.5187 0.4593 0.2581 0.6581 0.5223
BCE-R 0.5108 0.5314 0.4887 0.5777 0.4590 0.4788 0.4601 0.5414 0.5380 0.5095
AnomalyGFM 0.5974 0.8417 0.6751 0.9032 0.5791 0.5843 0.5380 0.6195 0.5614 0.6555
AUPRC FA 0.0400 0.0776 0.0960 0.0684 0.1555 0.2478 0.0356 0.1199 0.1495 0.1100
BCE 0.0394 0.0544 0.2254 0.0718 0.1594 0.2264 0.0272 0.0287 0.1389 0.1079
BCE-R 0.0314 0.0677 0.0685 0.0787 0.1294 0.2122 0.0305 0.0658 0.0971 0.0868
AnomalyGFM 0.0387 0.5790 0.1242 0.6820 0.1819 0.2749 0.0397 0.1371 0.0593 0.2352
We also evaluate the sensitivity of AnomalyGFM with subgraph sizes and the results are shown in Fig. 3. The performance of AnomalyGFM remains stable with varying subgraph sizes, demonstrating its robust generalization capability in modeling large graphs.
Figure 3. AnomalyGFM performance w.r.t subgraph size $s$.
4.5. Ablation Study
To verify the effectiveness of each component of AnomalyGFM, we designed three variants including (i) BCE which replaces the similarity measurement with the predicted probability from the BCE loss function for anomaly scoring. (ii) BCE Residual (BCE-R) which firstly removes the alignment loss and applies the BCE loss on the residual features to differentiate normal and abnormal nodes. Then, the predicted probability is used as the anomaly score. (iii) Feature Alignment (FA) which replaces the alignment on the residual feature with the alignment on the original feature. The results are shown in Table 5. FA achieves the highest AUROC on the Elliptic and T-Finance datasets, suggesting that learning prompts from the original features can also be effective as the original feature also maintains anomaly discriminability but they aren’t in the unified feature space leading to sub-optimal performance. BCE does not perform well except on the small dataset Disney. It directly uses the predicted probability as the anomaly score, which makes it struggle to generalize to new graphs due to the differences between the training and test sets. BCE-R does not achieve good performance compared to BCE. The reason is that although the residual features are discriminative, many samples located on the decision boundaries after being mapped to a unified feature space may not be effectively scored by the supervised classifier. In contrast, aligning the learnable prototype with the average node representation residual and using a similarity measurement for scoring can effectively discriminate the normal and abnormal nodes.
4.6. Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of AnomalyGFM w.r.t the size of the prototype $T$, hyperparameter $\alpha$, and commonality dimension $d^{\prime}$. Additional sensitivity analyses are in Appendix B.3.
AUROC and AUPRC results of AnomalyGFM w.r.t. prototype size $T$. As shown in Fig. 4, AnomalyGFM generally achieves better performance with increasing prototype size. This suggests that a larger prototype size provides more information and enhances the modeling capability of the foundation model.
Figure 4. AnomalyGFM performance w.r.t prototype size $T$.
AUROC and AUPRC results of AnomalyGFM w.r.t. $\alpha$. We vary the hyperparameter $\alpha$ in the range of [0.6, 1.2] with the interval of 0.1 and the results are shown in Fig. 5. We can see that AnomalyGFM remains stable with different $\alpha$ on most datasets in terms of both metrics, demonstrating the robustness of AnomalyGFM against this hyperparameter.
Figure 5. AnomalyGFM performance w.r.t $\alpha$.
Sensitivity of AnomalyGFM w.r.t common dimensionality $d^{\prime}$. To further demonstrate AnomalyGFM’s generalization, we evaluate AnomalyGFM under different $d^{\prime}$. The results are shown in Fig. 6. AnomalyGFM outperforms all the competing methods under different dimensions. The main reason is that the deviation between connected nodes is still preserved with different $d^{\prime}$ which can be efficiently measured by the graph-agnostic prototypes.
Figure 6. AUROC and AUPRC of AnomalyGFM w.r.t $d^{\prime}$
Impact of $\beta$ in anomaly scoring. The performance under different $\beta$ is shown in Fig. 7. As shown in Table 9 in App. B.3, the set of $\beta$ values are determined based on the prior we gain from the global average similarity. Specifically, the hyperparameter in anomaly scoring $\beta$ is set to zero by default for graphs with high global similarity such as Amazon and Yelp, and four for those with low global similarity such as Ellitipic and T-Finance, see the detailed global similarity information in Table 1. In the few-shot scoring, $\beta$ is set to $0.5$ by default for graphs with high global similarity. The main reason for these settings is that for graphs with high global average similarity, where node features are relatively similar, the deviations in local neighborhood would be small, and the fluctuations in scoring based on the normal prototype may affect the overall anomaly score. Therefore, we solely use the abnormal class prototype or assign a smaller weight for the normal class prototype in scoring. For nodes with less similar features, we use a mix of normal and abnormal prototypes for scoring. In the few-shot setting, the labeled normal nodes help refine the normal prototypes, allowing the anomaly scores to be computed using both normal and abnormal prototypes.
Figure 7. AUROC and AUPRC of AnomalyGFM w.r.t $\beta$.
Table 6. Training and inference time (seconds) comparison.
Methods AnomalyDAE TAM GAT BWGNN UNPrompt AnomalyGFM
Training Time 86.04 479.70 2.43 4.86 2.08 7.45
Inference Time 170.46 252.52 91.88 98.33 19.47 11.00
5. Time Complexity Analysis
In this section, we analyze the time complexity of AnomalyGFM. We build the GCN as the backbone of AnomalyGFM which takes $O(mdh)$, where $m$ is the number of non-zero elements in matrix $\bf A$ (the number of edge in graph), $d$ is the dimension of representation, and $h$ is the number of feature maps. The computation of the residual feature takes $O(N^{2}d)$. The MLP layer mapping the representation to the anomaly score takes $O(Nd)$. The feature mapping of normal class prototype and abnormal class prototype takes $O(2d^{2})$. The alignment loss between residual features and prototypes takes $O(2Nd)$. Thus, the overall complexity of the pre-training of AnomalyGFM is $O(mdh+N^{2}d+3Nd+2d^{2})$. In the inference, we employ the residual feature’s similarity with two prototypes which takes $O(2Nd)$. Therefore, the overall complexity of inference of AnomalyGFM is $O(mdh+N^{2}d+2d^{2}+2Nd)$. The complexity of inference on large-scale graph takes $O\left({K\left({mdh+{n^{2}}d+2{d^{2}}+2nd}\right)}\right)$ with $n$ as subgraph size and $K$ as the number of subgraphs. It is significantly lower than other methods, as they require loading entire graphs and performing reconstruction or affinity calculations, leading to higher space/time complexity. In Table 6, we report the running time comparison on Facebook and inference time on YelpChi-all. AnomalyGFM requires much less training and inference time compared to TAM and AnomalyDAE, demonstrating the efficiency. For inference, AnomalyGFM is the most efficient method because scoring relies solely on measuring the similarity between node representation residuals and prototypes.
6. Conclusion
In this paper, we build a GAD-oriented graph foundation model, AnoamlyGFM, that can work effectively under both few-shot and zero-shot scenarios. AnomalyGFM is pre-trained to learn discriminative and data-independent prototypes by aligning them with the graph-agnostic node representation residuals. This provides a consistent and identical way for abnormality measurement using the similarity between residual node representation and the learned class prototypes, facilitating the strong generalization in both zero-shot and few-shot inference. Extensive experiments on 11 datasets demonstrated the effectiveness and generalization of AnomalyGFM.
Acknowledgments
This research is supported by A*STAR under its MTC YIRG Grant (No. M24N8c0103), the Ministry of Education, Singapore under its Tier-1 Academic Research Fund (No. 24-SIS-SMU-008), the Lee Kong Chian Fellowship, and ARC under Grant DP240101349.
Appendix A Detailed Dataset Description
A detailed introduction of all the used datasets is given as follows.
  • Facebook (Xu et al., 2022): It is a social network where nodes represent users and edges indicate relationships between them. Anomalies are the nodes that either connect manually constructed clusters or display attributes differing from those of their contextual neighbors.
  • Reddit (Kumar et al., 2019): It is a user-subreddit graph, capturing one month’s worth of posts shared across various subreddits at Reddit. The users who have been banned by the platform are labeled anomalies. The text of each post is transformed into the feature vector and the features of the user and subreddits are the feature summation of the post they have posted.
  • Amazon (Dou et al., 2020): It includes product reviews under the Musical Instrument category. The users with more than 80% of helpful votes were labeled as begin entities, with the users with less than 20% of helpful votes treated as fraudulent entities.
  • Disney (Sánchez et al., 2013): It comes from the Amazon co-purchase network of movies where the attributes are the prices, ratings, number of reviews, etc. The anomalies are labeled manually by majority vote and the ground truths are derived from amazonfail stage information.
  • Amazon-all (Dou et al., 2020): It includes three types of relations: U-P-U (users reviewing at least one same product), U-S-U (users giving at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities). Amazon-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023).
  • YelpChi-all (Dou et al., 2020): Similar to Amazon-all, YelpChi-all includes three types of edges: R-U-R (reviews posted by the same user), R-S-R (reviews for the same product with the same star rating), and R-T-R (reviews for the same product posted in the same month). YelpChi-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023).
  • Tolokers (Platonov et al., 2023): It is obtained from Toloka crowdsourcing platform where the node represents the person who has participated in the selected project and the edge connects two workers means that they work on the same task. The attributes of the node are the profile and task performance statistics of workers.
  • Question (Platonov et al., 2023): It is collected from the website Yandex Q where the node represents the user and the edge connecting the node represents the user who has answered other’s questions during a one year period. The attribute is the mean of embeddings for words in the description of the user.
  • Elliptic (Weber et al., 2019): It is a Bitcoin transaction network in which the node represents the transactions and the edge is the flow of Bitcoin currency. The Bitcoin transaction is mapped to real-world entities associated with licit categories.
  • T-Finance (Tang et al., 2022): It is a financial transaction network where the node represents an account and the edge represents two accounts that have transaction. The features of each account are related to some attributes of logging like registration days, logging activities, and interaction frequency, etc. The users are labeled as anomalies when they fall into the categories like fraud money laundering and online gambling.
  • T-Social (Tang et al., 2022): It is a social network where the node represents the user and the edge indicates the relationship between users for more than three months. The attributes of each account are some logging information like registration days, logging activities, interaction frequency, etc.
Appendix B Additional Experimental Results
B.1. Few-shot Tuning on Abnormal Nodes
Additionally, tuning the abnormal prototype ${\bf{p}_{a}}$ is also feasible when few-shot abnormal node labels are available. Give some labeled abnormal nodes ${{\mathcal{V}}_{l}^{Test}}$, the fine tuning process is formulated as
(12) ${L_{{{pt}}}}=\sum\limits_{i=1}^{\left|{{\mathcal{V}}_{l}^{Test}}\right|}\left\|{{{\bf{r}}_{i}}-{\bf{p}^{\prime}_{a}}}\right\|_{2}^{2},$
where ${{\bf{p}^{\prime}_{a}}}={{\bf{p}}_{a}}+{g}({{\bf{p}}_{a};\phi})+\Psi_{\mathbf{p}_{a}}$, with a similar formulation applied during the tuning of few-shot normal nodes. To evaluate the performance of AnomalyGFM under few-shot abnormal nodes setting. we include a few labeled abnormal nodes, and treat the unlabeled nodes as normal. These can be combined to serve as the training source. We compare AnomalyGFM with the few-shot GAD method MetaGAD (Xu et al., 2024) and fully supervised methods BWGNN (Tang et al., 2022). We show results in Table 7 for the scenario where few-shot labeled abnormal nodes are available. As shown in Table 7, AnomalyGFM outperforms MetaGAD and BWGNN since they are limited by the scarcity of anomalies on the target graphs, whereas AnomalyGFM has better generalization since it only needs to adapt the pre-trained anomaly prototype with the few anomaly nodes to the target graphs.
Table 7. Comparison under 1/5-shot anomalies
Metric Setting Method Dataset
Amazon-all YelpChi-all
AUROC 1-shot BWGNN (ICML’22) 0.7441 0.5447
MetaGAD (DSAA’24) 0.8602 0.5434
AnomalyGFM 0.9155 0.5941
5-shot BWGNN (ICML’22) 0.8292 0.4759
MetaGAD (DSAA’24) 0.8558 0.5741
AnomalyGFM 0.9246 0.6004
AUPRC 1-shot BWGNN (ICML’22) 0.2947 0.1703
MetaGAD (DSAA’24) 0.7562 0.1681
AnomalyGFM 0.6956 0.1922
5-shot BWGNN (ICML’22) 0.4433 0.1408
MetaGAD (DSAA’24) 0.7567 0.1814
AnomalyGFM 0.7538 0.1935
B.2. Train the Model on Amazon and Evaluate on Other Datasets
To further demonstrate the effectiveness of AnomalyGFM, we pre-train the AnomalyGFM on Amazon and evaluate it on the other datasets. The AUROC and AUPRC results are shown in the Table 8. AnomalyGFM consistently outperforms competing methods from supervised methods, unsupervised methods, and generalist methods across most datasets in both AUROC and AUPRC, demonstrating its strong generalization when pre-trained on a different auxiliary dataset.
Table 8. Comparison of AUROC and AUPRC results on 8 real-world GAD datasets under zero-shot setting with the models trained on Amazon only.
Metric Method Dataset Avg. p-value
Facebook Reddit Disney YelpChi-all Tolokers Question Elliptic T-Finance
AUROC Unsupervised Methods
AnomalyDAE (ICASSP’20) 0.6123 0.5799 0.4938 0.4912 0.5080 0.5401 0.2769 0.2886 0.4738 0.0156
CoLA (TNNLS’21) 0.5427 0.4962 0.5455 0.4937 0.4680 0.4768 0.6654 0.4108 0.5123 0.0390
GADAM (ICLR’24) 0.6024 0.4720 0.3966 0.5289 0.5193 0.5575 0.2907 0.1463 0.4391 0.0234
Supervised Methods
BWGNN (ICML’22) 0.5441 0.4026 0.4196 0.5841 0.4430 0.5117 0.5471 0.6196 0.5089 0.0390
XGBGraph (NeurIPS’23) 0.4869 0.4869 0.4376 0.5869 0.4710 0.5430 0.5367 0.5950 0.5180 0.0781
Generalist Methods
GraphPrompt (WebConf’23) 0.3093 0.4511 0.7128 0.4994 0.5161 0.4720 0.6772 0.4753 0.5141 0.0546
UNPrompt (Arxiv’24) 0.7917 0.5356 0.6959 0.5448 0.4949 0.4467 0.3558 0.1805 0.5057 0.0078
AnomalyGFM 0.8279 0.6331 0.7135 0.5898 0.5305 0.5268 0.6010 0.5784 0.6251 /
AUPRC Unsupervised Methods
AnomalyDAE (ICASSP’20) 0.0675 0.0413 0.0583 0.1479 0.2253 0.0361 0.0630 0.0300 0.0836 0.0234
CoLA (TNNLS’21) 0.0468 0.0327 0.0717 0.1474 0.2038 0.0289 0.1427 0.0367 0.0888 0.0546
GADAM (ICLR’24) 0.0461 0.0299 0.0732 0.1602 0.2348 0.0424 0.0636 0.0255 0.0844 0.0156
Supervised Methods
BWGNN (ICML’22) 0.0289 0.0263 0.0494 0.1975 0.2073 0.0316 0.0992 0.0635 0.0879 0.0390
XGBGraph (NeurIPS’23) 0.0268 0.0315 0.0541 0.1994 0.1982 0.0426 0.0991 0.0538 0.0881 0.0546
Generalist Methods
GraphPrompt (WebConf’23) 0.0169 0.0298 0.1157 0.1481 0.2166 0.0425 0.1590 0.0464 0.0968 0.1484
UNPrompt (Arxiv’24) 0.2291 0.0340 0.0933 0.1767 0.2294 0.0274 0.0708 0.0258 0.1108 0.1953
AnomalyGFM 0.1094 0.0440 0.1593 0.1979 0.2505 0.0322 0.1112 0.0604 0.1206 /
B.3. Sensitivity Analysis
Impact of $\beta$ in Anomaly Scoring. As shown in Table 9, the set of $\beta$ values are determined based on the prior we gain from the global average similarity in Table 1. It is set to 0 by default for graphs with high global similarity and 4 for graphs with low global similarity. In the few-shot scoring, $\beta$ is set to $0.5$ by default for graphs with high global similarity and four for graphs with low global similarity, allowing the anomaly scores to be computed using both normal and abnormal class prototypes, where the normal class prototype is refined using the labeled normal nodes.
Table 9. The value of $\beta$ in anomaly scoring in both zero-shot and few-shot setting. Sim. indicate the global average edge similarity of the graph.
Setting Sim.¿0.5 Sim.¡=0.5
Zero-shot Setting 0 4
Few-shot Setting 0.5 4
Impact of $\mu$ and $\sigma$ in Gaussian distribution initialization. To evaluate the effectiveness of AnomalyGFM, we set different values for $\mu$ and $\sigma$ during the Gaussian distribution initialization for both the normal class prototype and the abnormal class prototype, see in the Table 10. The performance remains stable when we adjust $\mu$ and $\sigma$, indicating that AnomalyGFM is not dependent on the initial Gaussian distribution. This is mainly because the supervised loss function provides sufficient supervision, reducing the reliance on the Gaussian distribution.
Table 10. Performance of AnomalyGFM under different initialization hyperparameters for Gaussian distribution
Metric Initialization Dataset
Amazon-all Tolokers Question Elliptic T-Finance
AUROC $\mu$ = 0, $\sigma$ = 1 0.9032 0.5843 0.5280 0.6195 0.5614
$\mu$ = 0, $\sigma$ = 0.5 0.9229 0.5981 0.5378 0.6068 0.5902
$\mu$ = 0.1, $\sigma$ = 0.5 0.8756 0.6004 0.5437 0.5907 0.5727
$\mu$ = 0.2, $\sigma$ = 1 0.8963 0.5975 0.5429 0.6099 0.5922
AUPRC $\mu$ = 0, $\sigma$ = 1 0.6820 0.2749 0.0335 0.1333 0.0593
$\mu$ = 0, $\sigma$ = 0.5 0.7334 0.2824 0.0345 0.1234 0.0612
$\mu$ = 0.1, $\sigma$ = 0.5 0.6568 0.2870 0.0342 0.1096 0.0596
$\mu$ = 0.2, $\sigma$ = 1 0.6959 0.2853 0.0339 0.1177 0.0604
Appendix C Algorithms
The pre-training, zero-shot inference, few-shot prompt fine-tuning, and subgraph-based inference processes of AnomalyGFM are summarized in Algorithm 1, Algorithm 2, Algorithm 3, and Algorithm 4, respectively.
Algorithm 1 Pre-training of AnomalyGFM
1:  Input: Training graph $\mathcal{G}_{\rm{Train}}=({\mathcal{V}_{train},\mathcal{E}_{train}})$; Pre-training epoch $E$.
2:  Output: Pre-training $AnomalyGFM(\mathbf{A},\mathbf{X}).$
3:  Perform feature unification of $\mathbf{X}$ to obtain $\tilde{\mathbf{X}}$.
4:  Initialize the Gaussian distribution $\bf{z}_{n}$ and $\bf{z}_{a}$ for normal class prototype and abnormal class prototype.
5:  Randomly initialize GNN ${{(\mathbf{h}^{(0)}_{1},\mathbf{h}^{(0)}_{2},...,\mathbf{h}^{(0)}_{N})}\leftarrow{{\tilde{\bf{X}}}}}$
6:  for  ${epoch=1,\cdots,E}$  do
7:     for each ${v}$ in ${{\mathcal{V}_{train}}}$ do
8:        for  ${l=1,\cdots,L}$ do
9:           ${{\mathbf{h}}_{v}^{(l)}={{\mathbf{W}}^{(l)}\mathbf{h}}_{v}^{(l-1)}}$
10:           ${{\mathbf{h}}_{v}^{(l)}={\mathop{\rm R}\nolimits}{\rm{eLU}}\left({{\rm{AGG}}(\{{\mathbf{h}}_{v^{\prime}}^{(l)}:(v,v^{\prime})\in{\mathcal{E}}_{train}\})}\right)}$
11:        end for
12:     end for
13:     Compute the residual feature ${\bf{r}}_{i}$ of each node $v_{i}$ using Eq. (3).
14:     Compute the ${L_{BCE}}$ for classifier ${{p_{i}}={f_{\theta}}\left({\bf h}_{i}\right)}$.
15:     Alignment the normal class prototype and abnormal class prototype with the residual feature using ${L_{Alignment}}$
16:     Compute the total loss ${L_{{total}}}={L_{BCE}}+\beta{L_{Alignment}}$.
17:     Update the trainable weight parameters $\bf{W}$, ${\theta}$, $\Theta_{a}$ and $\Theta_{n}$ by using gradient descent.
18:  end for
19:  return  Pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$, $\Phi({\bf{z}}_{n};\Theta_{n})$ and $\Phi({\bf{z}}_{a};\Theta_{a}).$
Algorithm 2 Zero-shot Inference of AnomalyGFM
1:  Input: Testing graph $\mathcal{G}_{\text{test}}=({\mathcal{V}_{test},\mathcal{E}_{test}})$, pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$, $\Phi({\bf{z}}_{n};\Theta_{n})$ and $\Phi({\bf{z}}_{a};\Theta_{a})$.
2:  Output: Anomaly score of testing nodes.
3:  Initialization of the Gaussian distribution of normal class prototype and abnormal class prototype.
4:  for each $v_{i}$ in ${{\mathcal{V}_{test}}}$ do
5:     Apply the pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$ on the graph.
6:     Compute the residual feature ${\bf{r}}_{i}$ of each node $v_{i}$ using Eq. (3).
7:     Scoring using Similarity measurement with normal class prototype ${\bf p}_{n}$ and abnormal class prototype ${\bf p}_{a}$
8:     Compute the anomaly score using Eq. (9).
9:  end for
10:  return  Anomaly scores of testing nodes ${s(v_{1}),\cdots,s(v_{N})}$.
Algorithm 3 Few-shot Prompt Fine-tuning of AnomalyGFM
1:  Input: Testing graphs $\mathcal{G}_{\text{test}}=({\mathcal{V}_{test},\mathcal{E}_{test}})$, pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$, $\Phi({\bf{z}}_{n};\Theta_{n})$ and $\Phi({\bf{z}}_{a};\Theta_{a})$, the training epoch $e$ of fine-tuning.
2:  Output: The fine-tuned $AnomalyGFM(\mathbf{A},\mathbf{X})$ with updated normal class prototype ${\bf p}_{n}^{\prime}$.
3:  Initialization of the Gaussian distribution of normal class prototype and abnormal class prototype.
4:  Randomly initialize GNN ${{(\mathbf{h}^{(0)}_{1},\mathbf{h}^{(0)}_{2},...,\mathbf{h}^{(0)}_{N})}\leftarrow{{\tilde{\bf{X}}_{test}}}}$
5:  for  ${epoch=1,\cdots e}$  do
6:     for each ${v}$ in ${{\mathcal{V}_{test}^{l}}}$ do
7:        for  ${l=1,\cdots,L}$ do
8:           ${{\mathbf{h}}_{v}^{(l)}={{\mathbf{W}}^{(l)}\mathbf{h}}_{v}^{(l-1)}}$
9:           ${{\mathbf{h}}_{v}^{(l)}={\mathop{\rm R}\nolimits}{\rm{eLU}}\left({{\rm{AGG}}(\{{\mathbf{h}}_{v^{\prime}}^{(l)}:(v,v^{\prime})\in{\mathcal{E}}_{test}\})}\right)}$
10:        end for
11:        Compute the residual feature ${\bf r}_{i}$ of each labeled normal node $v_{i}$ using Eq. (3)
12:     end for
13:     Compute the $L_{pt}$ and optimize the newly added trainable parameters $\phi,\Psi$ using gradient descent.
14:  end for
15:  return  The fine-tuned $AnomalyGFM(\mathbf{A},\mathbf{X})$ with the updated normal class prototype ${\bf p}_{n}^{\prime}$.
Algorithm 4 Subgraph-based Inference of AnomalyGFM
1:  Input: Testing graphs $\mathcal{G}_{\text{test}}=({\mathcal{V}_{test},\mathcal{E}_{test}})$, pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$, $\Phi({\bf{z}}_{n};\Theta_{n})$ and $\Phi({\bf{z}}_{a};\Theta_{a})$.
2:  Output: Anomaly score of testing nodes.
3:  for each $v_{i}$ in ${{\mathcal{V}_{test}}}$ do
4:     Extract the subgraph ${\mathcal{S}}\left(v_{i}\right)$ of node $v_{i}$ using random walk.
5:     Apply the pre-trained $AnomalyGFM(\mathbf{A},\mathbf{X})$ on the extracted subgraph.
6:     Compute the residual feature ${\bf{r}}_{i}$ of each node $v_{i}$ using Eq. 11.
7:     Similarity measurement with normal class prototype ${\bf p}_{n}$ and abnormal class prototype ${\bf p}_{a}$.
8:     Compute the anomaly score using Eq. (9).
9:  end for
10:  return  Anomaly scores of testing nodes ${s(v_{1}),\cdots,s(v_{N})}$.